The logarithmic function f(x) = -100 + 87 ln x models the number of visitors (in millions) to a certain country's museums, where x is the number of years since the initial recording of the number of visitors. Use this function to estimate the number of visitors in year 60. Round to the nearest tenth of a year.
Answer by ankor@dixie-net.com(22740)

he logarithmic function f(x) = -100 + 87 ln x models the number of visitors (in millions) to a certain country's museums,
where x is the number of years since the initial recording of the number of visitors.
Use this function to estimate the number of visitors in year 60.
Round to the nearest tenth of a year ???
:
f(x) = -100 + 87*ln(60)
use your calc find ln(60)
f(x) = -100 + 87(4.0943)
f(x) = -100 + 356.208
f(x) = 256.208 ~ 256.2 million